OpenStudy (anonymous):

CALCULUS HELP!! Three forces of 5N, 8N and 10N act form the corner of a rectangular solid along its three edges. a. Calculate the magnitude of the equilibrate of these three forces. b. Determine the angle that the equilibrate makes with each of the three forces.

OpenStudy (dumbcow):

The resultant vector of these 3 forces is the sum of their components ---> R = <5,8,10> magnitude: \[|R| = \sqrt{25+64+100} = 3 \sqrt{21}\] the angles can be found using the relation: \[\cos \theta = \frac{u*v}{|u| |v|}\]